# One Model for All Tasks: Leveraging Efficient World Models in Multi-Task Planning

### v1

@@ Appendix A Implementation Details
This appendix details the implementation of our ScaleZero model, explorations of key design spaces, the Dynamic Parameter Scaling strategy, hyperparameter configurations, and the computational resources used for our experiments. Unless otherwise specified, the training and inference details remain consistent with those in the paper Pu et al. (2024).
A.1 Core Modifications of ScaleZero over UniZero
@@
Our core model, ScaleZero, is built upon the UniZero baseline, with key components upgraded to enhance learning efficiency and final performance in complex multi-task environments.
Visual Encoder: For image-based inputs (e.g., Atari), we employ a Vision Transformer (ViT) as the primary encoder. The input image (64x64x3) is divided into patches (8x8), which are then converted into token embeddings via a linear projection layer. Learnable positional embeddings are added, and the sequence is fed into a standard Transformer encoder (ViT-Base architecture) to produce a 768-dimensional latent state.
Text Encoder: In text-based environments like Jericho, we use a pre-trained BGE (BAAI General Embedding) model (Xiao et al., 2023) to encode the textual observations from the game into fixed-dimensional vectors.
Unified World Model:
  • Backbone: The core of the model is a Transformer network. It processes state tokens from the encoder, historical action tokens, and optional task-embedding tokens.
@@
  • UniZero (Baseline): The backbone consists of $N$ standard Transformer blocks. We use $N=8$ for the Atari26 benchmark and $N=4$ for other benchmarks.
@@
  • ScaleZero (Our Model): To maintain a comparable parameter count, the backbone is composed of $N/2$ Mixture-of-Experts (MoE) Transformer blocks. In each of these blocks, the standard feed-forward network (FFN) layer is replaced by an MoE layer. This layer comprises 8 expert networks and one shared expert, governed by a gating network that performs sparse routing.
Latent State Normalization: In ScaleZero, we apply standard LayerNorm to normalize the latent state.
A.2 Explorations on Key Design Spaces
@@
We conducted several exploratory extensions to key designs of UniZero to validate their effectiveness in multi-task settings.
@@
Task Information Encoding: To mitigate representational conflicts in multi-task learning, we explored explicit encoding of task information. By introducing a learnable task embedding matrix, task_embed = nn.Embedding(task_id), and concatenating it with the state representation, new_latent_state = concat(latent_state, task_embed), we aimed for the model to better distinguish between different tasks. In our experiments, the latent_state dimension was 672, and the task_embed dimension was 96.
@@
Multi-task Gradient Correction (MoCo): We adopted the MoCo algorithm from the LibMTL library [^3] https://github.com/median-research-group/LibMTL to handle gradient conflicts. Its core idea is to first maintain a momentum term $y_{t}$ to smoothly track each task’s gradient direction. Then, task weights $\lambda$ are dynamically updated by solving a regularized optimization problem. Finally, the gradient applied to shared parameters is the weighted sum of the momentum gradients. Key hyperparameters are listed in Table 5.
@@
Hyperparameter Value
MoCo_beta ($\beta$) 0.99
MoCo_gamma ($\gamma$) 10.0
MoCo_rho 0.0
A.3 Dynamic Parameter Scaling (DPS)
To address the issue of non-uniform learning progress across tasks in multi-task learning, we propose the Dynamic Parameter Scaling (DPS) strategy. The core idea is to dynamically align the model’s effective capacity with the set of currently unsolved tasks. This is achieved through a curriculum-based expansion of the model’s parameters. The detailed pseudocode is provided in Algorithm 1.
Staged Training and Active Set Management. The training protocol is divided into $S+1$ stages. We define a task $\tau_{i}$ as "solved" once its performance metric surpasses a predefined threshold $\varepsilon_{i}$. The set of all currently unsolved tasks forms the active set $\mathcal{U}_{t}$. To maximize efficiency, both data collection and training updates are performed exclusively on tasks within this active set.
  • Stage 0 (Warm-up): During an initial period of $T_{0}$ iterations, only the shared backbone network parameters $\theta_{B}$ are trained. This allows the model to learn a foundational, general-purpose representation across all tasks.
  • Stage $s\geq 1$ (Expansion): Each subsequent stage introduces new, dedicated parameters in the form of an independent Low-Rank Adaptation (LoRA) module, $\Delta\theta_{s}=B_{s}A_{s}$, and a corresponding learnable scalar scaling factor $\alpha_{s}$. These new parameters are intended to provide additional capacity for tackling the remaining difficult tasks in $\mathcal{U}_{t}$.
Adaptive Stage Transition Triggers. The transition from stage s1s-1 to stage $s$ is governed by a dual-trigger mechanism, ensuring a balance between progress and resource allocation:
  1. 1.
    Progress-Driven Trigger: A new stage is initiated if the number of newly solved tasks since the beginning of the current stage reaches a predefined quota $Q_{s}$. This allows the model to expand only when it has demonstrated mastery over a subset of the current challenges.
  2. 2.
    Budget-Driven Trigger: To prevent stagnation on particularly hard tasks, a transition is forced if the number of iterations within the current stage exceeds a maximum budget, calculated as $\lceil(T_{\max}-T_{0})/S\rceil$. This ensures that the model continues to gain new capacity even if progress stalls.
Hierarchical Parameter Composition and Optimization. In any given stage $s$, the forward pass for a weight matrix $W_{0}\in\theta_{B}$ is dynamically composed as:
$W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$ (3)
Upon entering a new stage $s$, we implement a crucial optimization strategy: we freeze the backbone parameters $\theta_{B}$ and all previously introduced LoRA modules $\{\Delta\theta_{j}\}_{j=1}^{s-1}$. The optimization gradient is then applied only to the newly introduced LoRA module $\Delta\theta_{s}=\{A_{s},B_{s}\}$ and to all existing scaling factors $\{\alpha_{j}\}_{j=1}^{s}$. This design provides two key benefits: it mitigates negative transfer by isolating new parameter updates within the fresh LoRA module, and it retains plasticity by allowing the model to dynamically re-weigh the contribution of knowledge from all previous stages via the trainable scalars $\{\alpha_{j}\}$.
The specific DPS settings used in our ScaleZero-Balanced experiments are detailed in Table 6.
Table 6: Hyperparameters for Dynamic Parameter Scaling (DPS).
Hyperparameter Value Description
curriculum_stage_num 5 Total number of expansion stages ($S$).
lora_r 64 The rank of the LoRA matrices ($r$).
lora_alpha 1 Initial scaling factor for LoRA layers.
min_stage0_iters 10,000 Duration of the warm-up stage ($T_{0}$).
max_stage_iters 5,000 Per-stage iteration budget for the budget-driven trigger.
@@ A.3 Dynamic Parameter Scaling (DPS)
A.4 Common Settings and Hyperparameters
To ensure a fair comparison, most hyperparameters are kept identical to the original UniZero paper. We use the AdamW optimizer. Key shared hyperparameters are summarized in Table 7.
Table 7: Common hyperparameters for all experiments.

### v2

@@ Appendix A Implementation Details
This section provides a comprehensive overview of the implementation details for ScaleZero and the associated experimental methodology. For clarity, we first establish a baseline by providing a concise summary of the UniZero baseline (Pu et al., 2024), including its training and MCTS procedures. Building upon this foundation, we then sequentially detail the core modifications that define ScaleZero, our explorations of key design spaces, the Dynamic Parameter Scaling (DPS) strategy, and finally, the full hyperparameter and computational setup to ensure reproducibility. Our code is available at https://github.com/opendilab/LightZero.
A.1 UniZero
As outlined above, we begin by detailing the foundational framework of our baseline, UniZero, to provide the necessary context for our contributions. The learning process revolves around two core procedures operating in a loop:
  1. 1.
    collect_experience: The agent interacts with the environment to gather trajectories. At each timestep, it uses Monte Carlo Tree Search (MCTS) within its learned latent world model to generate an improved policy, from which an action is sampled. The resulting experience tuples (observation, action, reward, done) and the MCTS-improved policy are stored in a replay buffer.
  2. 2.
    update_world_model: The world model, policy network, and value network are jointly optimized using batches of data sampled from the replay buffer. This step updates the model’s understanding of the environment’s dynamics and improves its decision-making capabilities.
A.1.1 MCTS in the Learned Latent Space
The MCTS procedure is central to UniZero’s planning capability and is used during the collect_experience phase. For each action selection, a search tree is built in the latent space over a fixed number of simulations. Each simulation consists of three phases:
  • Selection: Starting from the root node (the latent state of the current observation), the search traverses the tree by selecting actions that maximize an upper confidence bound score. This score, calculated using the PUCT formula, balances exploiting high-value actions (Q-value) and exploring less-visited actions (policy prior P). The action $a^{*}$ is chosen according to:
    $a^{*}=\arg\max_{a}\left[Q(\hat{z},a)+P(\hat{z},a)\frac{\sqrt{\sum_{b}N(\hat{z},b)}}{1+N(\hat{z},a)}\left(c_{1}+\log\left(\frac{\sum_{b}N(\hat{z},b)+c_{2}+1}{c_{2}}\right)\right)\right]$ (2)
    where $N$ is the visit count, $Q$ is the mean action-value, $P$ is the policy prior, and $c_{1},c_{2}$ are exploration constants.
  • Expansion: When a leaf node is reached, the dynamics network is called to predict the next latent state $\hat{z}^{l+1}$ and reward $\hat{r}^{l}$. The decision network predicts a policy $p^{l}$ and value $v^{l}$ for the new state. This new node is added to the tree, and its outgoing edges are initialized with the predicted policy priors.
  • Backup: The value $v^{l}$ predicted at the new leaf node is used to update the statistics of the nodes along the simulation path. The Q-values and visit counts $N$ for all parent state-action pairs are updated recursively back to the root.
After all simulations are complete, the visit counts at the root node are normalized to form a temperature-controlled distribution, which serves as the improved policy for action selection.
A.1.2 Model and Policy Optimization
The update_world_model step optimizes the network parameters by minimizing a joint loss function over batches of trajectories sampled from the replay buffer. The optimization objective in UniZero is composed of four main prediction losses:
$\begin{split}\mathcal{L}_{\text{UniZero}}(\theta)\doteq\mathop{\mathbb{E}}_{(\dots)\sim\mathcal{B}}\Big[\sum_{t=0}^{H-1}\Big(&\beta_{z}\underbrace{\|\hat{z}_{t+1}-\operatorname{sg}(\bar{h}(o_{t+1}))\|^{2}_{2}}_{\text{next-latent prediction}}+\beta_{r}\underbrace{\operatorname{CE}(\hat{r}_{t},r_{t})}_{\text{reward prediction}}\\ +&\beta_{p}\underbrace{\operatorname{CE}({p}_{t},\pi_{t})}_{\text{policy prediction}}+\beta_{v}\underbrace{\operatorname{CE}({v}_{t},\hat{v}^{targ}_{t})}_{\text{value prediction}}\Big)\Big]\end{split}$ (3)
where $H$ is the training context length and the loss components are defined as follows:
  • Next-latent Prediction: The model predicts the next latent state $\hat{z}_{t+1}$. The target is generated by a slowly-updating target encoder $\bar{h}$ and is detached from the computation graph using a stop-gradient operator ($\operatorname{sg}$).
  • Reward and Value Prediction: The model predicts the immediate reward $\hat{r}_{t}$ and the state value $v_{t}$. To handle varying scales across tasks, both predictions are framed as discrete classification problems and optimized using a cross-entropy (CE) loss. The value target $\hat{v}^{targ}_{t}$ is a bootstrapped $n$-step TD target computed using future rewards and a target value network.
  • Policy Prediction: The model’s policy head $p_{t}$ is trained to mimic the MCTS-improved policy $\pi_{t}$. This serves as a policy distillation step, which enhances training stability compared to direct policy gradient methods.
The terms $\beta_{z},\beta_{r},\beta_{p},\beta_{v}$ are fixed coefficients that balance the contribution of each component.
A.2 Core Modifications of ScaleZero over UniZero
@@
Our core model, ScaleZero, is built upon the UniZero baseline, with key components upgraded to enhance learning efficiency and final performance in complex multitask environments. The primary architectural differences are summarized in Table 3, with detailed explanations provided in the subsequent sections.
Table 3: Summary of core architectural differences between UniZero and ScaleZero.
Component UniZero (Baseline) ScaleZero (Our Model)
Backbone Standard Transformer blocks. Mixture-of-Experts (MoE) Transformer blocks.
Latent State Normalization SimNorm (L1 regularization). Standard LayerNorm.
Visual Encoder ResNet-like architecture trained from scratch. Vision Transformer (ViT) trained from scratch.
A.2.1 Backbone: From Dense Transformers to Sparse MoE
The most significant modification lies in the model’s backbone. While both models employ a Transformer architecture to process state, action, and task tokens, their internal structure differs fundamentally.
@@ A.2.1 Backbone: From Dense Transformers to Sparse MoE
  • UniZero (Baseline): The backbone consists of $N$ standard Transformer blocks (e.g., $N=8$ for Atari26).
@@
  • ScaleZero (Our Model): To maintain a comparable parameter count while increasing model capacity, the backbone is composed of $N/2$ Mixture-of-Experts (MoE) Transformer blocks. In each block, the standard feed-forward network (FFN) is replaced by an MoE layer. This layer comprises 8 specialized expert networks and one shared expert, governed by a sparse top-1 routing gate. This design creates conditional computation pathways, mitigating task interference.
A.2.2 Encoders and Normalization
Encoders
For encoding observations, ScaleZero adopts a similar strategy to the baseline but with a clear decision justified by empirical results.
  • Visual Encoder (e.g., Atari): Our primary encoder is a Vision Transformer (ViT) trained from scratch. The input image (64x64x3) is divided into 8x8 patches, projected into token embeddings, and processed by a standard ViT-Base encoder to produce a 768-dimensional latent state.
  • Text Encoder (e.g., Jericho): For text-based environments, we use a pre-trained BGE (BAAI General Embedding) model (Xiao et al., 2023) to encode textual observations into fixed-dimensional vectors.
Justification for From-Scratch Visual Encoder
In preliminary experiments, we explored leveraging a powerful pre-trained visual encoder, specifically a frozen DINOv2 model (dinov2_vits14) (Oquab et al., 2023), to potentially improve feature quality. However, this approach did not yield significant or consistent performance gains over the from-scratch ViT. We hypothesize this is due to the domain gap between the natural images DINOv2 was trained on and the distinct visual characteristics of Atari frames. This finding justifies our final choice and highlights the importance of domain-specific visual representations. We leave a more in-depth investigation of pre-trained models, including fine-tuning strategies, as future work.
Latent State Normalization
ScaleZero replaces the L1-based SimNorm used in UniZero with standard LayerNorm for normalizing the latent state. This change was found to be effective within our modified architecture.
A.2.3 Task-Specific Architectural Configurations
It is important to note that certain architectural choices are adapted based on the benchmark, rather than being a modification of ScaleZero over UniZero. These configurations apply to our entire framework.
  • Encoders: For the DeepMind Control (DMC) suite, which features heterogeneous vector observation spaces, we employ independent MLP encoders for each task. For benchmarks with a unified input modality (images for Atari, text for Jericho), a single shared encoder is used across all tasks.
  • Prediction Heads: To accommodate the distinct action spaces of each task, independent prediction heads (for policy, value, etc.) are universally employed across all benchmarks.
A.3 Explorations on Key Design Spaces
@@
We conducted several exploratory extensions to key designs of UniZero to validate their effectiveness in multitask settings.
@@
Task Information Encoding: To mitigate representational conflicts in multitask learning, we explored explicit encoding of task information. By introducing a learnable task embedding matrix, task_embed = nn.Embedding(task_id), and concatenating it with the state representation, new_latent_state = concat(latent_state, task_embed), we aimed for the model to better distinguish between different tasks. In our experiments, the latent_state dimension was 672, and the task_embed dimension was 96.
@@
Multitask Gradient Correction with MoCo. In multitask learning, gradients from different tasks often conflict, hindering overall progress. To mitigate this, we adapt the MoCo algorithm (Fernando et al., 2023), originally from the LibMTL library [^3] https://github.com/median-research-group/LibMTL, for our distributed training framework. The core idea is to correct the gradients applied to the shared model parameters by dynamically adjusting task weights based on smoothed gradient directions. This process involves two key steps: momentum-based gradient smoothing and dynamic weight adjustment.
  • (1). Momentum-based Gradient Smoothing. Instead of using the noisy, single-step gradient $g_{i}(t)$ for each task $i$ at step $t$, MoCo maintains a momentum-based moving average of the gradients, denoted as $y_{i}(t)$. This provides a more stable and reliable estimate of each task’s long-term descent direction. The update rule is:
    $y_{i}(t)=(1-\beta)y_{i}(t-1)+\beta g^{\prime}_{i}(t)$ (5)
    where $g^{\prime}_{i}(t)$ is the raw gradient $g_{i}(t)$ normalized and then scaled by its corresponding loss value, and $\beta$ is the momentum coefficient.
  • (2). Dynamic Weight Adjustment. The task weights, $\lambda(t)=\{\lambda_{1}(t),\dots,\lambda_{N}(t)\}$, are updated to minimize the squared norm of the aggregated gradient, i.e., $\min_{\lambda}\frac{1}{2}\|\sum_{i=1}^{N}\lambda_{i}y_{i}(t)\|^{2}$. Our implementation approximates a gradient descent step on this objective, followed by a softmax projection to ensure the weights form a valid probability distribution. The update for the weights is:
    $\lambda(t)=\text{softmax}\left(\lambda(t-1)-\gamma(Y(t)Y(t)^{T})\lambda(t-1)\right)$ (6)
    where $Y(t)$ is the matrix whose columns are the momentum gradients $\{y_{i}(t)\}$, and $\gamma$ is the learning rate for the weights. The final corrected gradient applied to the shared parameters $\theta_{s}$ is then the weighted sum $g_{s}(t)=\sum_{i=1}^{N}\lambda_{i}(t)y_{i}(t)$.
  • Our implementation is tailored for a Distributed Data Parallel (DDP) environment. All gradient correction computations occur on the rank-0 process, and the resulting corrected gradient is then broadcast to all other processes to ensure synchronized updates. The detailed procedure is outlined in Algorithm 1. Key hyperparameters are listed in Table 5.
Algorithm 1 MoCo in a Distributed Data Parallel (DDP) Setting
1:Shared parameters $\theta_{s}$, task losses $\{L_{i}\}_{i=1}^{N}$.
2:Momentum gradients $\{y_{i}\}_{i=1}^{N}$, task weights $\{\lambda_{i}\}_{i=1}^{N}$.
3:Hyperparameters: Momentum coefficient $\beta$, weight learning rate $\gamma$.
4:Updated shared parameters $\theta_{s}$.
5:for each training step do
[…]
7:  Compute local raw gradients $\{g_{i}^{\text{local}}\}$ for tasks on process pp.
8:  Gather all local gradients $\{g_{i}^{\text{local}}\}$ and losses $\{L_{i}^{\text{local}}\}$ to the rank-0 process.
[…]
10:  Let $\{g_{i}\}_{i=1}^{N}$ and $\{L_{i}\}_{i=1}^{N}$ be the full sets of gradients and losses.
11:  for $i=1$ to NN do
12:  Normalize and scale gradient: $g^{\prime}_{i}\leftarrow\frac{g_{i}}{\|g_{i}\|+\epsilon}\cdot L_{i}$.
13:  Update momentum gradient: $y_{i}\leftarrow(1-\beta)y_{i}+\beta g^{\prime}_{i}$.
14:  end for
15:  Let $Y\leftarrow[y_{1},\dots,y_{N}]$.
16:  Update task weights: $\lambda\leftarrow\text{softmax}(\lambda-\gamma(YY^{T})\lambda)$.
17:  Compute final corrected gradient: $g_{s}\leftarrow\sum_{i=1}^{N}\lambda_{i}y_{i}$.
18:  Broadcast the corrected gradient $g_{s}$ from rank-0 to all other processes.
[…]
20:  Set gradients of shared parameters $\nabla\theta_{s}\leftarrow g_{s}$.
21:  Perform optimizer step on $\theta_{s}$.
22:end for
@@
Hyperparameter Symbol Value
Momentum Coefficient $\beta$ 0.99
Weight Learning Rate $\gamma$ 10.0
Weight Regularization $\rho$ 0.0
A.4 Dynamic Parameter Scaling (DPS)
This section provides a detailed breakdown of the Dynamic Parameter Scaling (DPS) strategy introduced in Section 3.3. DPS is designed to dynamically couple model capacity with learning progress, creating a "curriculum of model complexity" that directs resources precisely where they are most required. Detailed pseudocode is provided in Algorithm 2.
@@ A.4 Dynamic Parameter Scaling (DPS)
Task Curation and Active Set Management. As described in the main text, we define the full set of tasks as $\mathcal{T}=\{\tau_{i}\}_{i=1}^{N}$. A task $\tau_{i}$ is deemed "solved" once its performance metric, $\text{Metric}(\tau_{i})$, surpasses a predefined threshold $\varepsilon_{i}$. At any time $t$ during training, we maintain an active set of all unsolved tasks, $\mathcal{U}_{t}\subseteq\mathcal{T}$. To maximize computational efficiency, both data collection and gradient updates are performed exclusively for tasks within this active set $\mathcal{U}_{t}$. Once a task is solved, it is removed from the active set, thereby ceasing all computational overhead associated with it.
Staged Capacity Expansion. The training process is partitioned into $S+1$ stages, following the protocol outlined in the main text:
  • Stage 0 (Warm-up): During an initial period of $T_{0}$ iterations, only the shared backbone network parameters, $\theta_{B}$, are trained. This stage allows the model to learn a foundational, general-purpose representation across all tasks.
  • Stage $s\geq 1$ (Expansion): Each subsequent stage introduces a new, independent LoRA module, $\Delta\theta_{s}$. This module comprises a trainable low-rank matrix pair $\Delta\theta_{s}=B_{s}A_{s}$ (with rank $r\ll d,k$) and a corresponding learnable scalar scaling factor $\alpha_{s}$. These new parameters provide additional model capacity for tackling the remaining difficult tasks in the active set.
Adaptive Stage Transition Triggers. The transition from stage s1s-1 to stage $s$ is governed by a dual-trigger mechanism, balancing learning progress against a fixed budget:
  1. 1.
    Progress-based Trigger: A new stage is initiated if the number of newly solved tasks within the current stage reaches a predefined quota $Q_{s}$.
  2. 2.
    Budget-based Trigger: To prevent stagnation on particularly arduous tasks, a transition is forced if the iteration count within the current stage exceeds a pre-allocated limit, calculated as $\lceil(T_{\max}-T_{0})/S\rceil$.
Optimization and Parameter Isolation. In any given stage $s$, the forward pass for a weight matrix $W_{0}\in\theta_{B}$ is dynamically composed as:
$W^{(s)}(x)=W_{0}x+\sum_{j=1}^{s}\alpha_{j}(B_{j}A_{j})x$ (7)
Crucially, upon entering a new stage $s$, all prior parameters—the backbone $\theta_{B}$ and the matrices of all previously introduced adapters $\{A_{j},B_{j}\}_{j=1}^{s-1}$—are frozen. Optimization is thereby focused exclusively on the newly added LoRA module $\Delta\theta_{s}=\{A_{s},B_{s}\}$ and on the full set of currently active scaling factors $\{\alpha_{j}\}_{j=1}^{s}$. In practice, this is achieved by using ‘torch.detach()‘ on the outputs of adapters from previous stages, which stops their gradients while preserving their forward contribution.
This strategy provides two decisive advantages:
  • It implements a resource-efficient training curriculum, as new parameters are only introduced when required by the remaining unsolved tasks.
  • By isolating new learning within dedicated adapters, DPS provides targeted plasticity for difficult tasks while preventing catastrophic forgetting or negative transfer to previously mastered skills, whose knowledge is preserved in the frozen backbone and prior adapters.
Stabilized Learnable Scaling Factors. To prevent training instability from unbounded learnable scalars, we constrain the scaling factors $\alpha_{j}$ within a narrow, stable range (e.g., $[0.8,1.2]$). This is achieved via a re-parameterization trick, where each $\alpha_{j}$ is computed from an unbounded underlying parameter $\hat{\alpha}_{j}$ as follows:
$\alpha_{j}=\text{offset}+\text{range}\cdot\tanh(\hat{\alpha}_{j})$ (8)
This design ensures that the contribution of each adapter is modulated smoothly without causing drastic shifts in the output distribution. The specific DPS hyperparameters used in our ScaleZero-DPS experiments are detailed in Table 6.
Table 6: Hyperparameter configuration for Dynamic Parameter Scaling in our ScaleZero-DPS experiments.
Hyperparameter Symbol Value Description
curriculum_stage_num $S+1$ 5 Total number of stages (1 warm-up + 4 expansion).
lora_r $r$ 64 The rank of the LoRA matrices.
lora_alpha - 1 Initial value for LoRA scaling (for ‘scaling = alpha / r‘).
lora_scale_init $\alpha_{j,\text{init}}$ 1.0 Initial value of the learnable scaling factors.
lora_scale_range - 0.2 Range for learnable scalars, yielding $\alpha_{j}\in[0.8,1.2]$.
min_stage0_iters $T_{0}$ 10,000 Number of iterations for the warm-up stage.
max_stage_iters - 5,000 Per-stage iteration budget, i.e., $\lceil(T_{\max}-T_{0})/S\rceil$.
A.5 Hyperparameter Settings
Our hyperparameter configuration for ScaleZero largely follows the original UniZero paper to ensure a fair comparison. We employ the AdamW optimizer for all experiments. Table 7 lists the hyperparameters that are kept consistent across all benchmarks. Architectural configurations and loss weights are detailed subsequently, followed by benchmark-specific settings in Tables 9, 10, and 11.
Table 7: Hyperparameters shared across all benchmarks.